# Evaluating and Improving the Robustness of LiDAR-based Localization and Mapping

### v1

@@ IV Experiment Setup
IV-A Subject LiDAR SLAM Systems
@@
SLAM MULLS [76] 2021 Feature-based 0.56%
@@
DeLORA [78] 2021 Projection-based 7.78%
NeRF-LOAM [79] 2023 Distance Field 1.70%
Map-based
Localization LocNDF [37] 2023 Distance Field 0.059m
We experiment four SOTA LiDAR SLAM techniques and one map-based LiDAR localization, as shown in Table IV. We try to select one representative work from different types of mainstream LiDAR localization technologies, i.e. feature-based, direct, and projection-based methods [4]. We also include two neural distance field-based approaches, which have emerged and become popular since the success of the NeRF [80] in computer vision.
MULLS [76] is a feature-based SLAM that uses ground filtering and principal components analysis to extract features from the PCD. It builds local sub-maps using previous PCD and match the current frame with local sub-maps by the proposed multi-metric linear least square iterative closest point algorithm to compute the pose. KISS-ICP [77] is a direct SLAM that combines point-to-point ICP with adaptive thresholding, a robust kernel, universal motion compensation, and point cloud subsampling. DeLORA [78] is a projection-based, unsupervised SLAM. It projects the 3D point cloud to 2D image representations and employs ResNet to extract features from the 2D image. Then it predicts the translation and rotation by feeding the features into an MLP. NeRF-LOAM employs a neural network to predict the signed distance field (SDF). It first builds the previous frames into an octree. Then it samples intersected points between laser rays and voxels in the previously built octree. Given sampled points, the neural network predicts the SDF. Since sampled points are transformed to the same coordinate system as the previously built octree by the pose to be estimated, NeRF-LOAM optimizes the pose, neural network, and voxel embedding jointly by minimizing the loss between predicted SDF and real SDF. Different from NeRF-LOAM, LocNDF trains a neural network offline to predict the SDF. Given the online LiDAR scan, it tries to find a pose $P$ so that the neural network can produce minimal SDF when sampled points are transformed by $P$.
@@ IV-B Datasets
We use KITTI Visual Odometry/SLAM Evaluation 2012 (KITTI) [19] to evaluate the four SLAM systems, DeLORA, MULLS, KISS-ICP, and NeRF-LOAM. KITTI 00 to 10 sequences are released with ground truths. We conducted our experiment on sequences 09 and 10 because (1) DeLORA was trained on sequences 00-08 hence only 09 and 10 remain. (2) sequences 09 and 10 are of moderate sizes which facilitate the large number of perturbations in a reasonable amount of time and computation resources.
@@ IV-C Evaluation Metrics
For the evaluation of the precision of SLAMs, we adopted the relative pose error (RPE) [19]. RPE is widely used in previous work [4] to measure and compare the performance of SLAM. RPE is calculated as:
@@ V Experiment Results
V-A Robustness Results of LiDAR SLAM
Method. We applied our robustness framework to evaluate the subject SLAM systems (Table IV). We first perturb the KITTI and Apollo-SouthBay datasets with the corruptions in Table I. Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24GB GPU memory and days to finish KITTI for one of the 18 perturbations), we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
@@ V-A Robustness Results of LiDAR SLAM
For the non-learning SLAM systems, we find that they have degraded performance on almost all kinds of LiDAR perturbed data compared to the baseline (i.e., un-perturbed data). Furthermore, each SLAM system is vulnerable to a particular type of perturbation, where their performance drops significantly. The error of MULLS increases from below 0.5% to up to 0.8%. Even worse, the error of MULLS with “local density decrease” (severity 5) increased to more than 12% for KITTI 09 and 18% for KITTI 10 from 0.5%, which is more than 100 times worse. Similarly, kiss-icp is especially vulnerable to “background noise”. While the error only increased by less than 0.2% in most cases, the error of kiss-icp caused by background noise increases up to 80% (with severity 5 on KITTI 10) from less than 0.6% RPEtrans. Moreover, even the lightest background noise (severity 1) causes a significant error increase from less than 0.6% to more than 3% RPEtrans on KITTI 10.
@@
We observe different results for the learning-based SLAM, Delora and NeRF-LOAM. Delora is vulnerable to most perturbations, especially those related to noise. The RPEtrans error increases from less than 10 percent to tens of percent (up to 50%). Lastly, NeRF-LOAM is quite different from others. We find the results differ very much between the two KITTI sequences, indicating that NeRF-LOAM’s robustness results highly depend on data. In KITTI 09, NeRF-LOAM is not sensitive to most perturbations except for one. While in KITTI 10, NeRF-LOAM is shown to be highly sensitive in most cases. Also, interestingly, in many cases, it shows better robustness with a higher severity than a lower severity. For example, for the cutout perturbation on KITTI 10, NeRF-LOAM performs poorly with severity 4, however, the performance is very well and much higher with severity 5 which is close to the baseline with the clean data. NeRF-LOAM is optimizing a mini-network online for tens of iterations. We hypothesize that the poor results are because NeRF-LOAM is far from being optimized in tens of iterations for these cases.
@@
For the map-based localization system LocNDF, our evaluation shows that it is very robust against all types of corruptions. The map-based SLAM LocNDF is robust, at the cost of significant upfront investments (in terms of multiple scans of the environment to build the LiDAR map) and its inability to work in unexplored environments.
@@
Our results show that SOTA Map-based localization, LocNDF, is robust against LiDAR data corruption under adverse weather and noises. The other SOTA SLAM systems including two learning-based SLAM are vulnerable to one or multiple types of LiDAR data corruption.
V-B Improving the Robustness of LiDAR SLAM
@@
Method. We only concern MULLS, kiss-icp, and Delora in this denoising experiment, since LocNDF is robust against all kinds of corruption, while NeRF-LOAM could not produce consistent results between different trials on the same data. Firstly, we apply the Bilateral Filter to denoise the corrupted data and then re-run all the localization algorithms on the denoised data. Furthermore, since Delora uses a learning-based approach, we augment Delora’s training set (i.e., KITTI 00 to 08 sequences) by applying our proposed LiDAR corruptions and re-train Delora on the augmented data.
@@ Results.
We present the performance of the three systems on the denoised data for KITTI 09 in Figure 4 (a). The results on KITTI 10 are available in our replication package due to the space limitation. For MULLS, local density decrease (local_dec) at severity 5 is the corruption that affects it most significantly. We find that the errors even increased by 2.43% after denoising. Besides, we find that denoising could reduce the error by 0.26% for Gaussian noise. The changes are insignificant for the remaining cases (less than 0.1%).
@@
Bilateral Filter is beneficial to noise-based corruption, and can fully eliminate the negative effects of background noise. However, it does not help with density-based corruption (local density decrease).
@@
Re-training with the corruption data is a more effective way to improve the robustness of the learning-based SLAM (Delora).
@@
In this work, we extensively evaluated the robustness of the LiDAR-based localization systems, including SLAM and map-based localization. Our results revealed that all kinds of SOTA LiDAR-based SLAM systems are vulnerable to specific types of LiDAR data corruption. We further showed that simple Bilateral Filter denoising can effectively eliminate noise-based data corruptions, however, the density-based corruptions remain unsolved. We also showed that the data augmentation with corrupted data can improve the robustness of the learning-based LiDAR SLAM significantly, and is even beneficial to its performance on clean data.
VII Acknowledgement
We would like to express our gratitude to our interns, Marko Griesser-Aleksic, Ali Ouabou El Ghoulb, and Hanchong Qiu, for their invaluable assistance and contribution to this project.

### v2

@@ IV Experiment Setup
IV-A Subject LiDAR-Based Pose Estimation Systems
@@
Odometry MULLS [24] 2021 Feature-based 0.56%
@@
Delora [19] 2021 Learning-based 7.78%
NeRF-LOAM [22] 2023 NeRF-based 1.70%
Localization LocNDF [21] 2023 NeRF-based 0.059m
We experiment with four SOTA LiDAR odometry and one LiDAR localization systems, as shown in Table IV. We cover one representative work from different types of mainstream point cloud registration methods, i.e., feature-based, direct, and learning-based methods [10]. We also include two neural distance field-based approaches, which have become popular since the success of the NeRF [20].
MULLS [24] is a feature-based LiDAR odometry that uses ground filtering and principal components analysis to extract features from the point cloud. It builds local sub-maps using the previous point cloud and match the current frame with local sub-maps by the proposed multi-metric linear least square iterative closest point algorithm to compute the pose. KISS-ICP [15] is a direct matching LiDAR odometry that combines point-to-point ICP with adaptive thresholding, a robust kernel, universal motion compensation, and point cloud subsampling. Delora [19] is a learning-based, unsupervised LiDAR odometry. It projects the 3D point cloud to 2D image representations and employs ResNet to extract features from the 2D image. Then it predicts the translation and rotation by feeding the features into a neural network. NeRF-LOAM [22] employs a neural network to predict the signed distance field (SDF). It first builds the previous frames into an octree. Then it samples intersected points between laser rays and voxels in the previously built octree. Given sampled points, the neural network predicts the SDF. Since sampled points are transformed to the same coordinate system as the previously built octree by the pose to be estimated, NeRF-LOAM optimizes the pose, neural network, and voxel embedding jointly by minimizing the loss between predicted SDF and real SDF. Different from NeRF-LOAM, LocNDF trains a neural network offline to predict the SDF. Given the online LiDAR scan, it tries to find a pose $P$ so that the neural network can produce minimal SDF when sampled points are transformed by $P$.
@@ IV-B Datasets
We use KITTI Visual Odometry/SLAM Evaluation 2012 (KITTI) [41] to evaluate the four odometry systems, Delora, MULLS, KISS-ICP, and NeRF-LOAM. KITTI 00 to 10 sequences are released with ground truths. We conducted our experiment on sequences 09 and 10 because (1) Delora was trained on sequences 00-08 hence only 09 and 10 remain. (2) sequences 09 and 10 are of moderate sizes which facilitate the large number of perturbations in a reasonable amount of time and computation resources.
@@ IV-C Evaluation Metrics
For the evaluation of the precision of pose estimation, we adopted the relative pose error (RPE) [41]. RPE is widely used in previous work [10] to measure and compare the performance of pose estimation. RPE is calculated as:
@@ V Experiment Results
V-A Evaluating Robustness of LiDAR-Based Pose Estimation
Method. We applied our robustness framework to evaluate the subject odometry systems. We first perturb the KITTI and Apollo-SouthBay datasets with the corruptions in Table I. Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24GB GPU memory and days to finish one a single sequence), so we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
@@ V-A Evaluating Robustness of LiDAR-Based Pose Estimation
For the non-learning systems, we find that they have degraded performance on almost all kinds of LiDAR perturbed data compared to the baseline (i.e., un-perturbed data). Furthermore, each system is vulnerable to a particular type of perturbation, where their performance drops significantly. The error of MULLS increases from below 0.5% to up to 0.8% on most of the corruptions. However, the error of MULLS with “local density decrease” (severity 5) increased to more than 12% for KITTI 09 and 18% for KITTI 10. Similarly, kiss-icp is especially vulnerable to “background noise”. While the error only increased by less than 0.2% in most cases, the error of kiss-icp caused by background noise increases up to 80% (with severity 5 on KITTI 10) from less than 0.6% RPEtrans. Moreover, even the lightest background noise (severity 1) causes a significant error increase from less than 0.6% to more than 3% RPEtrans on KITTI 10.
@@
We observe different results for the learning-based systems, Delora and NeRF-LOAM. Delora is vulnerable to most perturbations, especially those related to noise. The RPEtrans error increases from less than 10 percent to tens of percent (up to 50%). Lastly, NeRF-LOAM is quite different from others. We find the results differ very much between the two KITTI sequences, indicating that NeRF-LOAM’s robustness results highly depend on data. In KITTI 09, NeRF-LOAM is not sensitive to most perturbations except for one. While in KITTI 10, NeRF-LOAM is shown to be highly sensitive in most cases. Also, interestingly, in many cases, it shows better robustness with a higher severity than a lower severity. For example, for the cutout perturbation on KITTI 10, NeRF-LOAM performs poorly with severity 4, however, the performance is very well with severity 5 which is close to the baseline. NeRF-LOAM optimizes a mini-network online for tens of iterations. We hypothesize that the poor and unstable results are because NeRF-LOAM is far from being optimized in tens of iterations for these cases.
@@
For the LiDAR localization system LocNDF, our evaluation shows that it is very robust against all types of corruption. The robustness is at the cost of significant upfront investments (in terms of multiple scans of the environment to build the LiDAR map) and its inability to work in unexplored environments.
@@
Our results show that the SOTA LiDAR localization system, LocNDF, is robust against all point cloud corruptions. SOTA LiDAR odometry systems are vulnerable to one or multiple types of point cloud corruptions.
V-B Improving Robustness of LiDAR-Based Pose Estimation
@@
Method. We only concern MULLS, kiss-icp, and Delora in this denoising experiment, since LocNDF is robust against all kinds of corruption, while NeRF-LOAM could not produce consistent results between different trials on the same data. Firstly, we apply the bilateral filter to denoise the corrupted data and then re-run all the algorithms on the denoised data. Furthermore, since Delora uses a learning-based approach, we augment Delora’s training set (i.e., KITTI 00 to 08 sequences) by applying the proposed point cloud corruptions and re-train Delora on the augmented data.
@@ Results.
We present the performance of the three systems on the denoised data for KITTI 09 in Figure 3 (a). The results on KITTI 10 are similar and available in our replication package. For MULLS, local density decrease (local_dec) at severity 5 is the corruption that affects it most significantly. We find that the errors even increased by 2.43% after denoising. Besides, we find that denoising could reduce the error by 0.26% for Gaussian noise. The changes are insignificant for the remaining cases (less than 0.1%).
@@
Bilateral filter can effectively mitigate noise-induced corruption in general, and fully eliminate the negative effects of background noise. However, it does not help with density-related corruption (local density decrease).
@@
Re-training with the corruption data is a more effective way to improve the robustness of the learning-based LiDAR pose estimation system (Delora).
@@
In this work, we extensively evaluated the robustness of LiDAR-based pose estimation systems, including odometry and localization, against 18 synthetic real-world point cloud corruptions. Our findings indicate that all SOTA LiDAR-based odometry systems are particularly vulnerable to certain corruption types. We further show that simple bilateral filtering effectively mitigates noise-induced data corruptions, however, density-related corruptions remain challenging. Finally, we show that corruption-based data augmentation not only significantly enhances the robustness of learning-based LiDAR odometry systems but also improves their performance on clean data.